0.1g of a metal gave on reaction with dil. acid at STP 34.2ml hydrogen gas. Th equivalent weight of the metal is ???

Grade:12th pass

3 Answers

Vikas TU
7 years ago
Moles at STP => V/22.4 => 34.2/22400 = > 1.526 millimoles.
Eq. weight = 0.1/  1.526 x 10^-3 => 65.5307 gm
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
Moles at STP => V/22.4 => 34.2/22400 = > 1.526 millimoles.
Eq. weight = 0.1/  1.526 x 10^-3 => 65.5307 gm
 
Thanks and regards,
Kushagra
Ramya
2 years ago
Equivalent weight of metal=equivalent weight of hyrogen
0.1/equivalent weight of metal=34.2*2/22400
Therefore equivalent weight of metal=32.7